# ARES: Alternating Reinforcement Learning and Supervised Fine-Tuning for Enhanced Multi-Modal Chain-of-Thought Reasoning Through Diverse AI Feedback

### v1

@@ Appendix A Prompts
| Feedback | Prompt Structure |
| Correction Feedback |
[Prompt when Image is provided]
@@ Appendix B Difficulties with External Knowledge in A-OKVQA
Appendix C Training Details
C.1 Reinforcement Learning
Common Settings: We use top- sampling with and and sample 4 actions. The initial coefficient for the Kullback-Leibler (KL) divergence is set to . The range for clipping the probability ratios in PPO is . The discount factor is set to . Token length is constrained to . We train the model using 4 NVIDIA A100 80GB GPUs.
C.2 Supervised Fine-Tuning
C.3 LoRA Adapter Training
Appendix D Comparison of Generated Rationales
@@
| Look at each object. For each object, decide if it has that property. A sticky object can attach or stick to other things. The rubber duck is not sticky. Yellow is a color. This color is yellow. The rubber duck and the rain boots are yellow, but the lemon is not.nA rough object feels scratchy when you touch it. All three objects are rough. The property that all three objects have in common is rough. | |

### v2

@@ 6 Conclusion
Acknowledgments
We sincerely appreciate the generous support of computational resources provided by the Ohio Supercomputer Center (Center, 1987).
@@ Appendix B Difficulties with External Knowledge in A-OKVQA
Appendix C Training Details
C.1 Reinforcement Learning
Common Settings: We use top- sampling with and sample 4 actions. The initial coefficient for the Kullback-Leibler (KL) divergence is set to . The range for clipping the probability ratios in PPO is . The discount factor is set to . Token length is constrained to . We train the model using 4 NVIDIA A100 80GB GPUs.
C.2 Supervised Fine-Tuning
C.3 LoRA Adapter Training
C.4 Time Comparison between LoRA Adapter and Inference Model
Rather than introducing a separate model for inference, we achieve comparable performance by adding the LoRA adapter to the rationale model, while simultaneously obtaining a 2x–14x speedup in inference (ARES) compared to MM-CoT, which introduces a separate inference model. We provide the time comparison in Table 7 and Table 8.
This speed gap is mainly due to the fact that the separate inference model requires rationale generation before the inference procedure. However, the LoRA adapter directly refers to the rationale model’s latent features to derive answers, eliminating the need to generate the rationale first.
Model Rationale Inference Total 1h 32m 7m 1h 39m - 8m 8m 5h 23m 12m 5h 35m - 24m 24m
Model Rationale Inference Total 6m 2m 8m - 3m 3m 16m 3m 19m - 6m 6m
Appendix D Comparison of Generated Rationales
| Feedback | Prompt Structure |
| Correction Feedback |
[Prompt when Image is provided]
@@ Appendix D Comparison of Generated Rationales
| Comparison of Corrected Rationales After RL | |
| Image |
| Question | Will these magnets attract or repel each other? |
| Choices | repel, attract |
| After RL | Will these magnets attract or repel? To find out, look at which poles are closest to each other. The north pole of one magnet is closest to the south pole of the other magnet. Poles that are the same repel. So, these magnets will repel each other. The picture shows two magnets that are the same shape and the same size. Both magnets are the same size. So, these magnets will repel each other. |
| Corrected Rationale | The magnets will attract each other because the north pole of one magnet is closest to the south pole of the other magnet. Opposite poles of magnets attract each other in this diagram. Therefore, the correct answer is that the magnets will attract. |
| Question | How long is a pen? |
| Choices | 20 kilometers, 20 centimeters |
| After RL | The better estimate for the length of a pen is 20 centimeters. 20 kilometers is too long. 20 centimeters is too short. 20 kilometers is too long. |
| Corrected Rationale | The better estimate for the length of a pen is 20 centimeters. 20 kilometers is an unreasonable estimate, as it is far too long for a typical pen. 20 centimeters is a more accurate and reasonable estimate for the length of a standard pen. |
| Image |
| Question | What is the capital of South Carolina? |
| Choices | Columbia, Montgomery, Charleston, Harrisburg |
| After RL | Columbia is the capital of South Carolina.nThe city of Columbia is Charleston, South Carolina. Charleston is the smallest city in South Carolina.nHayeburg is the smallest city in South Carolina. Montgomery is the smallest city in South Carolina. But the city of Columbia is home to many cultural and historical landmarks. |
| Corrected Rationale | Columbia is the correct answer as it is the capital city of South Carolina. It is the largest city in the state and serves as the seat of government. The other options, Montgomery (the capital of Alabama), Charleston (a city in South Carolina), and Harrisburg (the capital of Pennsylvania), are not the capital of South Carolina. |
@@
| Look at each object. For each object, decide if it has that property. A sticky object can attach or stick to other things. The rubber duck is not sticky. Yellow is a color. This color is yellow. The rubber duck and the rain boots are yellow, but the lemon is not. A rough object feels scratchy when you touch it. All three objects are rough. The property that all three objects have in common is rough. | |